# MLLMs Need 3D-Aware Representation Supervision for Scene Understanding

### v1

@@ A.1 World Coordinate Computation
Given a set of $N$ images =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$, each image $I_{i}$ is paired with its depth map $D_{i}\in\mathbb{R}^{H\times W}$, camera intrinsic matrix $K_{i}\in\mathbb{R}^{3\times 3}$, and camera-to-world extrinsic matrix $T_{i}\in\mathbb{R}^{4\times 4}$. For a pixel at $(u,v)$ in image $I_{i}$, the corresponding 3D coordinate in the global coordinate system, denoted as $\mathbf{C}_{i}(u,v)\in\mathbb{R}^{3}$, is computed as:
@@ A.2 Datsests for Training
For model fine-tuning, we utilize a collection of well-established 3D vision-language datasets. Specifically, we follow the model finetuning settings of Video-3D LLM video3dllm by using the validation splits of ScanRefer, Multi3DRefer, Scan2Cap, and ScanQA, as well as the test split of SQA3D. Across these datasets, the number of data samples varies significantly: ScanRefer and Scan2Cap each provide 36,665 samples, while Multi3DRefer offers 43,838 entries. ScanQA contains 26,515 instances, and SQA3D is the largest with 79,445 samples. Most datasets are derived from 562 unique scans, except SQA3D, which includes 518 scans. We further report the average lengths of questions and answers for each dataset. For example, question lengths range from approximately 13 to 38 words, with Scan2Cap and ScanQA also providing answer texts, averaging 17.9 and 2.4 words in length, respectively. In SQA3D, the average question and answer lengths are 37.8 and 1.1 words.
@@ A.3 Detailed Comparison
Scanrefer. Tab. 6 shows that our method 3DRS achieves the best overall performance on the ScanRefer validation set, especially in the challenging “Multiple” scenario where precise target discrimination is required. These results demonstrate that 3DRS effectively leverages multi-view images for robust spatial understanding and accurate object localization.
Multi3DRefer. In Tab. 7, 3DRS achieves the best overall results on the Multi3DRefer validation set, with top F1 scores in both standard and challenging scenarios. Our method consistently outperforms previous approaches, especially in the difficult zero-target and distractor settings, demonstrating superior robustness and spatial understanding.
@@
ScanQA. In Tab. 8, 3DRS achieves the best performance on the ScanQA validation set across almost all metrics, including EM, BLEU scores, METEOR, and CIDEr, demonstrating its strong effectiveness for 3D question answering.
@@
SQA3D. In Tab. 10, 3DRS achieves the highest scores on the SQA3D test set, outperforming all previous approaches on almost every question type as well as in the overall average, which demonstrates its superior capability for 3D question answering across diverse scenarios.
Scan2cap. In Tab. 10, 3DRS achieves the best performance on the Scan2Cap validation set in terms of CIDEr (C), and remains highly competitive on other metrics such as BLEU-4, METEOR, and ROUGE-L, demonstrating strong overall effectiveness for 3D captioning.
@@ C B-4 M R
A.4 Qualitative Results
Figs. 5 and 6 provide a visual summary of how our method performs on three challenging 3D scene understanding tasks. These tasks include identifying objects based on language, generating descriptions for specific regions, and answering spatial questions about the scene.
@@ A.5 Broader Impacts
A.6 Limitation
While our paper aims to enhance the 3D-awareness of MLLMs, the relatively limited size of the dataset used for finetuning—especially when compared to that used during the MLLM pretraining stage—may restrict the full realization of our approach’s potential. Consequently, the improvements demonstrated in this work may only represent an initial step toward more robust 3D understanding. A promising direction for future research is to incorporate 3D-awareness learning into the pretraining stage of MLLMs, which could lead to fundamentally stronger models with deeper 3D comprehension.

### v2

@@ A.1 World Coordinate Computation
Given a set of $N$ images $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{N}\}$, each image $I_{i}$ is paired with its depth map $D_{i}\in\mathbb{R}^{H\times W}$, camera intrinsic matrix $K_{i}\in\mathbb{R}^{3\times 3}$, and camera-to-world extrinsic matrix $T_{i}\in\mathbb{R}^{4\times 4}$. For a pixel at $(u,v)$ in image $I_{i}$, the corresponding 3D coordinate in the global coordinate system, denoted as $\mathbf{C}_{i}(u,v)\in\mathbb{R}^{3}$, is computed as:
@@ A.2 Datsests for Training
For model fine-tuning, we utilize a collection of well-established 3D vision-language datasets. Specifically, we follow the model finetuning settings of Video-3D LLM video3dllm by using the validation splits of ScanRefer, Multi3DRefer, Scan2Cap, and ScanQA, as well as the test split of SQA3D. Across these datasets, the number of data samples varies significantly: ScanRefer and Scan2Cap each provide 36,665 samples, while Multi3DRefer offers 43,838 entries. ScanQA contains 26,515 instances, and SQA3D is the largest with 79,445 samples. Most datasets are derived from 562 unique scans, except SQA3D, which includes 518 scans. We further report the average lengths of questions and answers for each dataset. For example, question lengths range from approximately 13 to 38 words, with Scan2Cap and ScanQA also providing answer texts, averaging 17.9 and 2.4 words in length, respectively. In SQA3D, the average question and answer lengths are 37.8 and 1.1 words. For the evaluation on VSI-Bench, we use the pre-training data from VG-LLM vg-llm.
@@ A.3 Detailed Comparison
Scanrefer. Tab.˜7 shows that our method 3DRS achieves the best overall performance on the ScanRefer validation set, especially in the challenging “Multiple” scenario where precise target discrimination is required. These results demonstrate that 3DRS effectively leverages multi-view images for robust spatial understanding and accurate object localization.
Multi3DRefer. In Tab.˜8, 3DRS achieves the best overall results on the Multi3DRefer validation set, with top F1 scores in both standard and challenging scenarios. Our method consistently outperforms previous approaches, especially in the difficult zero-target and distractor settings, demonstrating superior robustness and spatial understanding.
@@
ScanQA. In Tab.˜9, 3DRS achieves the best performance on the ScanQA validation set across almost all metrics, including EM, BLEU scores, METEOR, and CIDEr, demonstrating its strong effectiveness for 3D question answering.
@@
SQA3D. In Tab.˜11, 3DRS achieves the highest scores on the SQA3D test set, outperforming all previous approaches on almost every question type as well as in the overall average, which demonstrates its superior capability for 3D question answering across diverse scenarios.
Scan2cap. In Tab.˜11, 3DRS achieves the best performance on the Scan2Cap validation set in terms of CIDEr (C), and remains highly competitive on other metrics such as BLEU-4, METEOR, and ROUGE-L, demonstrating strong overall effectiveness for 3D captioning.
@@ C B-4 M R
A.4 Ablation Study
Table 12: Distillation on multiple layers.
Supervision ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 C@0.5 B-4@0.5 C EM EM
Last layer 62.9 56.1 60.4 54.9 41.6 86.1 104.8 30.3 60.6
Last layer + last 3rd layer 61.5 54.8 60.1 54.9 41.4 84.4 101.4 29.2 60.5
Last layer + last 3rd + last 5th layer 60.5 53.9 59.0 53.8 40.0 81.1 102.9 30.0 59.6
Table 13: Performance with different distillation losses.
Supervision ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 B-4@0.5 C@0.5 C EM EM
Euclidean loss 62.9 56.1 60.4 54.9 41.6 86.1 104.8 30.3 60.6
Cosine loss 62.2 55.5 60.4 55.2 41.8 85.9 104.5 30.1 60.7
Cosine + Euclidean 62.3 55.7 60.3 55.0 42.1 85.8 102.7 29.7 60.2
Table˜12 shows the effect of applying supervision to different numbers of network layers across multiple 3D scene understanding tasks, including object localization, captioning, and question answering. Supervising only the last layer consistently achieves the best performance on all benchmarks. As more intermediate layers are added for supervision, the results degrade. This suggests that multi-layer supervision may over-constrain geometric features and weaken semantic representations, ultimately hindering downstream performance. Future work may explore more advanced strategies to balance geometric and semantic cues.
Table˜13 reports the impact of different distillation loss functions, including euclidean loss, cosine loss, and their combination, across various 3D scene understanding benchmarks. The results show that all loss types yield very similar performance, indicating that the choice of feature distance metric has limited influence in our setting.
A.5 Qualitative Results
Visualizations Fig.˜4 illustrates qualitative results of our method across three tasks: visual grounding, object captioning, and question answering.
Figs.˜5 and 6 provide a visual summary of how our method performs on three challenging 3D scene understanding tasks. These tasks include identifying objects based on language, generating descriptions for specific regions, and answering spatial questions about the scene.